Project L requires an initial outlay at t = 0 of $35,000, its expected cash inflows are $8,000 per year for 9 years, and its WACC is 12%. What is the project's NPV? Do not round intermediate calculations. Round your answer to the nearest cent.

Project L requires an initial outlay at t = 0 of $57,569, its expected cash inflows are $11,000 per year for 8 years, and its WACC is 10%. What is the project's IRR? Round your answer to two decimal places.

Project L requires an initial outlay at t = 0 of $54,000, its expected cash inflows are $9,000 per year for 10 years, and its WACC is 14%. What is the project's payback? Round your answer to two decimal places.

Travis Industries plans to issue perpetual preferred stock with an $11.00 dividend. The stock is currently selling for $90.50, but flotation costs will be 10% of the market price, so the net price will be $81.45 per share. What is the cost of the preferred stock, including flotation? Round your answer to two decimal places.
Answer #1

1)

NPV = Present value of cash inflows - present value of cash outflows

NPV = Annuity * [1 - 1 / (1 + r)n] / r - Initial investment

NPV = 8,000 * [1 - 1 / (1 + 0.12)9] / 0.12 - 35,000

NPV = 8,000 * [1 - 0.36061] / 0.12 - 35,000

NPV = 8,000 * 5.32825 - 35,000

NPV = $7,626.00

2)

IRR is the rate of return that makes initial investment equal to present value of cash inflows

Initial investment = Annuity * [1 - 1 / (1 + r)n] / r

57,569 = 11,000 * [1 - 1 / (1 + r)8] / r

Using trial and error method i.e., after trying various values for R, lets try R as 10.53%

57,569 = 11,000 * [1 - 1 / (1 + 0.1053)8] / 0.1053

57,569 = 11,000 * [1 - 0.448909] / 0.1053

57,569 = 11,000 * 5.23353

57,569 = 57,569

Therefore, IRR is 10.53%

3)

Project's payback = Initial cost / cash flows

Project's payback = 54,000 / 9,000

Project's payback = 6.00 years

4)

Cost of preferred stock = (Preferred dividend / price after flotation cost) * 100

Cost of preferred stock = (11 / 81.45) * 100

Cost of preferred stock = 13.51%